OpenStudy (anonymous):

MEDAL & FAN ANYBODY ? A quadratic function f(x) has two complex roots: x = 3 – i and x = 3 + i. Which one of the following is the correct expression for f(x)? a. f(x) = x2 + 9 b. f(x) = x2 – 6x + 10 c. f(x) = x2 – 6x + 9 d. f(x) = x2 – 6x – 9

OpenStudy (freckles):

if x=3-i then x-(3-i) is a factor if x=3+i then x-(3+i) is a factor I multiply (x-(3-i))(x-(3+i))

OpenStudy (anonymous):

Is it D? @freckles

OpenStudy (freckles):

we can multiply the above to find out

OpenStudy (freckles):

have you tried that

OpenStudy (anonymous):

Its b

OpenStudy (freckles):

\[(x-(3-i))(x-(3+i)) \\ x \cdot x+x \cdot -(3+i)-(3-i) \cdot x +(3-i)(3+i)\] yes that sounds a lot better :)